need help on polymer chemistry question

Determine which polymerization mechanism(s), i.e. radical, anionic, and cationic, can be used for the following monomers. Briefly explain your answer.

H2C= CHCEH H2C= C(CH3)2 H3CHC=CHCH
Answer #1

Solution:

A) CH2=CHC6H5 (styrene)

Styrene undergo anionic polymerization to produce polystyrene. The presence of strong electron withdrawing (phenyl) group decreases the electron density and hence increases the ease of anionic polymerization.

B) CH2 = C(CH3)2

The given alkene is unsymmetrical, hence presence of electron donating group (methyl) increases the ease of cationic polymerization.

C) CH3-CH = CH-CH3

The given alkene is a symmetrical alkene. Therefore, the homolytic cleavage more prefered and formation of free radical is greatly increased. Thus, mechanism follow free radical polymerization.